### v1

@@ 3.2 Effective graphs
From 4, a limit state $\bm{X}^{*}$ in our topic-weighted HK model is an absorbing state. Similarly, from 5, a limit state $\bm{X}^{*}$ in our topic-weighted DW model is almost surely an absorbing state. If $\bm{X}^{*}$ is an absorbing state, then for each edge $(i,j)$ in $G$, we have either (1) $d_{\mathrm{max}}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})=0$, and therefore, the edge $(i,j)$ is in $G_{\mathrm{eff}}^{*}$; or (2) $d_{\mathrm{max}}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})\geq c$, and therefore the edge $(i,j)$ is not in $G_{\mathrm{eff}}^{*}$. Recall that for each distinct limit opinion, the set of nodes that have the same limit opinion is a limit opinion cluster. Therefore, the nodes that make up a connected component of $G_{\mathrm{eff}}^{*}$ belong to the same limit opinion cluster. However, there may not be a one-to-one correspondence between limit opinion clusters and connected components of $G_{\mathrm{eff}}^{*}$. In particular, if two or more connected components of $G_{\mathrm{eff}}^{*}$ have the same corresponding limit opinion, then they all correspond to the same limit opinion cluster.
@@
Let $\bm{X}(t)$ be the opinion state of our topic-weighted DW model with update rule (8) and initial opinion state $\bm{X}(0)$ and define $A_{\mathrm{eff}}(t)$ and $A_{\mathrm{eff}}^{*}$ as in (33) and (34). Then it almost surely holds that
@@
From Theorem 5, we have for all edges $(i,j)\in E$, the maximum discordance $d_{\max}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})$ is either 00 or greater than or equal to $c$. We prove this theorem by considering three cases for edge $(i,j)\in E$: (1) $d_{\max}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})=0$; (2) $d_{\max}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})>c$; and (3) $d_{\max}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})=c$.
@@
[…]
where $\{t_{s_{r}}\}_{r=1}^{\infty}\subset\{t_{s}\}_{s=1}^{\infty}$ is a subsequence (note that $\lim_{r\rightarrow\infty}t_{s_{r}}=\infty$). Because $i$ and $j$ interact on topic $k$ with a positive probability, we choose the pair $(i,j)$ to interact on the topic $k$ infinitely often. We use a similar idea as in Figure 1 to show that this scenario almost never occurs. For $\varepsilon=c\mu/2(1-\mu)$, there exists a time $T\geq 0$ such that $|x_{i}^{k}(t)-x_{i}^{*k}|<\varepsilon$ for all nodes $i$, all topics $k$, and all times $t\geq T$. With probability $1$, we choose the pair $(i,j)$ and the topic $k$ at some time $t\geq T$ in the subsequence $\{t_{s_{r}}\}_{r=1}^{\infty}$. When this occurs, by (42), we update the opinions using (8) and obtain
@@ 3.3 Region of receptiveness
which, for a node $i$ with opinion $\bm{x}$, is the region of opinions that a node $j$ can have so that $i$ and $j$ are receptive to each other on all topics. We omit the dependence of the parameters of the BCMs (such as $\omega$ and $c$) in our definition to simplify the notation.
@@
Let $\bm{u}_{\textrm{tr}}=(u^{1},\ldots,u^{s})^{T}$ denote a truncated vector of the first $s$ topics. We can rewrite (56) as the matrix equation $(\omega I+\bm{v}\bm{w}^{T})\bm{u}_{\mathrm{tr}}=c\bm{1}$, where $\bm{v}=\frac{1-\omega}{K}\bm{1}$, $\bm{w}=\bm{1}$, and $\bm{1}$ is $s$-dimensional vector whose entries are all $1$. By the Sherman–Morrison formula[2, 27], the matrix $\omega I+\bm{v}\bm{w}^{T}$ is invertible with its inverse equal to
@@
We now show that for $\omega=0$, the vertices of the closure of $\mathcal{U}$ are given by $\mathcal{V}_{0}\cup\mathcal{V}_{1}$. We observe that when $\omega=0$, equation (52) reduces to the same equation for all $k$. Therefore, to have a system of $K$ distinct equations, we have at most $1$ equation of the form (52) with the rest taking the form (53). If a system of equations has exactly one equation of the form (52), then the solution has $K-1$ zero entries and only one nonzero entry whose value is $cK$. If a system of equations has no equation of the form (52), then the entries of the solution are all 00. In both cases, the solutions are feasible. Therefore, the set of vertices of the closure of $\mathcal{U}$ is $\mathcal{V}_{0}\cup\mathcal{V}_{1}$.
@@
Define $\tilde{\mathcal{U}}=\{\bm{\sigma}\odot\bm{u}:\bm{\sigma}\in\{1,-1\}^{K},\bm{u}\in\mathcal{U}\}$, where $\odot$ denotes the entry-wise vector product and the entries of $\bm{\sigma}$ are either $1$ or $-1$. In other words, we obtain $\tilde{\bm{u}}\in\tilde{\mathcal{U}}$ by flipping the sign of one or multiple entries of $\bm{u}\in\mathcal{U}$. Using this notation, we have $\bm{y}\in\mathcal{R}_{\bm{x}}$ if and only if there is some $\tilde{\bm{u}}\in\tilde{\mathcal{U}}$ such that $\bm{y}=\bm{x}+\tilde{\bm{u}}$. Because $\mathcal{R}_{\bm{x}}$ is a translation of $\tilde{\mathcal{U}}$, it is enough to show that $\tilde{\mathcal{U}}$ is the interior of a bounded convex polytope. We note that $\tilde{\bm{u}}\in\tilde{\mathcal{U}}$ if and only if $\tilde{\bm{u}}$ satisfies
@@
for all topics $k=1,\ldots,K$ and all subsets $S\subseteq\{1,\ldots,K\}$. We also note that $\overline{\tilde{\mathcal{U}}}$ is bounded because the points it comprises satisfy $\|\bm{u}\|_{d,k}<c$ (see Footnote 7). Therefore, $\tilde{\mathcal{U}}$ is the interior of a bounded convex polytope, and its vertices satisfy a linear system comprised of equations of the form
@@
(64) $[\bm{1}_{S}]^{k}=\begin{cases}-1&k\in S\,,\\ 1&\textrm{otherwise}\,.\end{cases}$
@@
Suppose $\bm{u}=\bm{1}_{S}\odot\bm{v}$, where $S\subseteq\{1,\ldots,K\}$ and $\bm{v}$ is a vertex of $\mathcal{U}$. By Lemma 3.9, we must have $\bm{v}\in\mathcal{V}_{s}$ for $s\in\{0,1,\ldots,K\}$. We claim that $\bm{u}$ is a vertex if and only if $s>0$. Suppose $s>0$. We note that $\bm{v}$ must be the solution of a linear system with $s$ equations of the form (52) and $K-s$ equations of the form (53). Without loss of generality, we can assume that $\bm{v}$ is a solution of the linear system consisting of (52) for $k=1,\ldots,s$ and (53) for $k=s+1,\ldots,K$. This implies that $\bm{u}$ is a solution to the system consisting of equations (63) for $k=1,\ldots,s$ and (53) for $k=s+1,\ldots,K$. However, it is readily checked that this system is equivalent to the system consisting of equations (63) for $k=1,\ldots,s$ and equations
@@
In Fig. 2, we show the region of receptiveness in 2D for a fixed value of $c$ and various values of $\omega$. When $\omega=1$ (i.e., as in the baseline BCMs), the region of receptiveness is a square. As we decrease $\omega$, the area of the region of receptiveness increases; a node has a larger region of opinion space that a neighbor can be in for them to be mutually receptive on a particular topic.
@@
For the case $K=3$, the region of receptiveness for the opinion vector $\bm{x}$ consists of a polygon with 26 vertices of the form $\bm{x}+\bm{u}$. The $s=3$ case yields 8 vertices, where each entry of $\bm{u}$ is $c$ or $-c$. The $s=2$ case yields 12 vertices, where one entry of $\bm{u}$ is 00 and the remaining two entries are ${3c}/({\omega+2})$ or $-{3c}/({\omega+2})$. For the case $s=1$, it yields 6 vertices, where two entries of $\bm{u}$ are 00 and the remaining entry is ${3c}/({2\omega+1})$ or $-{3c}/({2\omega+1})$.
@@ 4 Details of our numerical simulations
In this section, we discuss the setup for our numerical simulations of our topic-weighted BCMs. For our simulations, we consider opinions with $K=2$ topics that lie in the region $[0,1]\times[0,1]$. For this bounded opinion space, we correspondingly examine confidence bounds888 When $c=0$, adjacent nodes $i$ and $j$ are never receptive to each other on any topic $k$. When $c=1$, adjacent nodes $i$ and $j$ are always receptive to each other on any topic $k$. We do not examine these values of $c$. $c\in(0,1)$.
@@ 4.2 Characterizing opinion fragmentation in our simulations
Researchers have introduced various notions of opinion polarization and fragmentation [10] and have developed multiple methods to quantify them [10, 16, 1]. Intuitively, one simple way to quantify opinion fragmentation is to look at the number of opinion clusters, with a larger number of opinion clusters indicating more opinion fragmentation. However, this does not take into account the cluster sizes (i.e., the number of nodes in a cluster). Suppose that there are two opinion clusters. If the two opinion clusters are of the same size, then one can view the opinions in the system as being more fragmented than if one opinion cluster has a large majority of the nodes and the other opinion cluster has a small minority. In our numerical simulations, there can be considerable variation in the sizes of the opinion clusters. Therefore, it is important to consider both the numbers of opinion clusters and their sizes to quantify opinion fragmentation in our numerical simulations.
To quantify opinion fragmentation, we use a modification of the order parameter $Q$ defined by Wang et al. [53]. We calculate [^10] Wang et al. [53] used the inequality $|x_{i}-x_{j}|\leq c$ in their calculation of $Q$ for the standard HK model. However, we use a strict inequality to be consistent with the strict inequality in the update rules (7) and (8). the order parameter $Q$ as
@@
where $\bm{1}_{d_{\mathrm{max}}(\bm{x}_{i}(t),\bm{x}_{j}(t))<c}$ is an indicator function that equals 1 when $d_{\mathrm{max}}(\bm{x}_{i}(t),\bm{x}_{j}(t))<c$ and equals 00 otherwise. The order parameter $Q$ indicates the fraction of pairs of nodes that are receptive to each other. When $Q=1$, all nodes are receptive to each other and we say that nodes reach a consensus state in the infinite-time limit, which is an opinion state in which all opinion vectors are equal. When $Q<1$, we say that the nodes are in a fragmented state. We say that smaller values of $Q$ indicate more opinion fragmentation. For a complete graph, the order parameter gives the fraction of edges $(i,j)\in E$ that are in the effective graph. In a graph that is not complete, some nodes are not adjacent to each other, so the order parameter is more difficult to relate to the effective graph.
@@
For each of our simulations, we characterize opinions by calculating the final order parameter $Q(T_{f})$ at the time at which our simulation reaches the stopping criterion (see section 4.1). Let $Q^{*}$ denote the limit order parameter, which is the order parameter (see (70)) calculated with the limit opinion vectors $\bm{x}_{i}^{*}$ for each node $i$. If $Q(T_{f})=1$, then there is a single final opinion cluster and all pairs of adjacent nodes are receptive to each other on all topics. In this scenario, nodes will compromise their opinions over time and eventually reach a consensus state, yielding a limit order parameter $Q^{*}=1$. If $Q(T_{f})<1$, then there are at least two final opinion clusters. In this scenario, the final opinion clusters need not be the same as the opinion clusters of $G_{\mathrm{eff}}^{*}$ (see discussion in section 4.1), and it is possible that $Q(T_{f})\neq Q^{*}$. However, for a small tolerance value (see discussion in section 4.1 ), the value of $Q(T_{f})$ is a reasonable approximation of the $Q^{*}$.
@@ 4.3 Initial opinion distributions
if $x^{1},x^{2}\in[0,1]$ and 00 otherwise, where $\delta$ is the Dirac delta function. We show this wedge distribution in Fig. 3. In this wedge distribution, the marginal opinion distributions for $x^{1}$ and $x^{2}$ (the opinions on topics 1 and 2, respectively) are uniform on $[0,1]$, which is identical to the independent case. However, $x^{1}$ and $x^{2}$ are highly correlated. To sample from this wedge distribution $f(x^{1},x^{2})$, we draw $x^{1}$ first uniformly at random from $[0,1]$ and then compute $x^{2}$ using $x^{2}=1-|2x^{1}-1|$.
@@
if $x^{1},x^{2}\in[0,1]$ and 00 otherwise. Here, $x^{1}$ and $x^{2}$ are the opinions on topics 1 and 2, respectively, $\sigma$ is the standard deviation of the opinions on topics 1 and 2, and $\rho$ is the Pearson correlation between the opinions on topics 1 and 2. In particular, we consider two truncated Gaussian distributions in our simulations: (1) $G(0.22,0)$, where $x^{1}$ and $x^{2}$ have the same variance and are independent; (2) $G(0.22,0.8)$, where $x^{1}$ and $x^{2}$ have the same variance but are correlated. To draw opinions randomly from a truncated Gaussian distribution, we first draw randomly from a non-truncated Gaussian distribution. If the sample does not lie in $[0,1]\times[0,1]$, we reject it and resample using the same procedure. For our choice of $\sigma=0.22$, we expect that 95% of the random samples lie within the opinion space $[0,1]\times[0,1]$, which makes this simple rejection sampling method effective.

### v2

@@ 3.2 Effective graphs
From 4, a limit state $\bm{X}^{*}$ in our topic-weighted HK model is an absorbing state. Similarly, from 5, a limit state $\bm{X}^{*}$ in our topic-weighted DW model is almost surely an absorbing state. If $\bm{X}^{*}$ is an absorbing state, then for each edge $(i,j)$ in $G$, we have either (1) $d_{\mathrm{max}}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})=0$, and therefore, the edge $(i,j)$ is in $G_{\mathrm{eff}}^{*}$; or (2) $d_{\mathrm{max}}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})\geq c$, and therefore the edge $(i,j)$ is not in $G_{\mathrm{eff}}^{*}$. Recall that for each distinct limit opinion, the set of nodes that have that limit opinion is a limit opinion cluster. Therefore, the nodes that make up a connected component of $G_{\mathrm{eff}}^{*}$ belong to the same limit opinion cluster. However, there may not be a one-to-one correspondence between limit opinion clusters and connected components of $G_{\mathrm{eff}}^{*}$. In particular, if two or more connected components of $G_{\mathrm{eff}}^{*}$ have the same corresponding limit opinion, then they all correspond to the same limit opinion cluster.
@@
Let $\bm{X}(t)$ be the opinion state of our topic-weighted DW model with update rule (8) and initial opinion state $\bm{X}(0)$, and define $A_{\mathrm{eff}}(t)$ and $A_{\mathrm{eff}}^{*}$ as in (33) and (34). Then it almost surely holds that
@@
From Theorem 5, we have for all edges $(i,j)\in E$, the maximum discordance $d_{\max}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})$ is either 0 or greater than or equal to $c$. We prove this theorem by considering three cases for edge $(i,j)\in E$: (1) $d_{\max}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})=0$; (2) $d_{\max}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})>c$; and (3) $d_{\max}(\bm{x}_{i}^{*},\bm{x}_{j}^{*})=c$.
@@
(42) $d_{k}(\bm{x}_{i}(t_{s_{r}}),\bm{x}_{j}(t_{s_{r}}))<c\text{~for all times~}t_{s_{r}}\,$,
where $\{t_{s_{r}}\}_{r=1}^{\infty}\subset\{t_{s}\}_{s=1}^{\infty}$ is a subsequence (note that $\lim_{r\rightarrow\infty}t_{s_{r}}=\infty$). Because $i$ and $j$ interact on topic $k$ with a positive probability, we choose the pair $(i,j)$ to interact on the topic $k$ infinitely often. We use a similar idea as in Figure 1 to show that this scenario almost never occurs. For $\varepsilon=c\mu/2(1-\mu)$, there exists a time $T\geq 0$ such that $|x_{i}^{k}(t)-x_{i}^{*k}|<\varepsilon$ for all nodes $i$, all topics $k$, and all times $t\geq T$. With probability 11, we choose the pair $(i,j)$ and the topic $k$ at some time $t\geq T$ in the subsequence $\{t_{s_{r}}\}_{r=1}^{\infty}$. When this occurs, by (42), we update the opinions using (8) and obtain
@@ 3.3 Region of receptiveness
which, for a node $i$ with opinion $\bm{x}$, is the region of opinions that a node $j$ can have so that $i$ and $j$ are receptive to each other on all topics. We omit the dependence of the parameters of the BCMs (namely, $c$ and $\omega$) in our definition to simplify the notation.
@@
Let $\bm{u}_{\textrm{tr}}=(u^{1},\ldots,u^{s})^{T}$ denote a truncated vector of the first $s$ topics. We can rewrite (56) as the matrix equation $(\omega I+\bm{v}\bm{w}^{T})\bm{u}_{\mathrm{tr}}=c\bm{1}$, where $\bm{v}=\frac{1-\omega}{K}\bm{1}$, $\bm{w}=\bm{1}$, and $\bm{1}$ is an $s$-dimensional vector whose entries are all $1$. By the Sherman–Morrison formula[3, 28], the matrix $\omega I+\bm{v}\bm{w}^{T}$ is invertible with its inverse equal to
@@
We now show that for $\omega=0$, the vertices of the closure of $\mathcal{U}$ are given by $\mathcal{V}_{0}\cup\mathcal{V}_{1}$. We observe that when $\omega=0$, equation (52) reduces to the same equation for all $k$. Therefore, to have a system of $K$ distinct equations, we have at most $1$ equation of the form (52) with the rest taking the form (53). If a system of equations has exactly one equation of the form (52), then the solution has K1K-1 zero entries and only one nonzero entry whose value is $cK$. If a system of equations has no equation of the form (52), then the entries of the solution are all 0. In both cases, the solutions are feasible. Therefore, the set of vertices of the closure of $\mathcal{U}$ is $\mathcal{V}_{0}\cup\mathcal{V}_{1}$.
@@
Define $\tilde{\mathcal{U}}=\{\bm{\sigma}\odot\bm{u}:\bm{\sigma}\in\{1,-1\}^{K},\bm{u}\in\mathcal{U}\}$, where $\odot$ denotes the entry-wise vector product and the entries of $\bm{\sigma}$ are either $1$ or 1-1. In other words, we obtain $\tilde{\bm{u}}\in\tilde{\mathcal{U}}$ by flipping the sign of one or multiple entries of $\bm{u}\in\mathcal{U}$. Using this notation, we have $\bm{y}\in\mathcal{R}_{\bm{x}}$ if and only if there is some $\tilde{\bm{u}}\in\tilde{\mathcal{U}}$ such that $\bm{y}=\bm{x}+\tilde{\bm{u}}$. Because $\mathcal{R}_{\bm{x}}$ is a translation of $\tilde{\mathcal{U}}$, it is enough to show that $\tilde{\mathcal{U}}$ is the interior of a bounded convex polytope. We note that $\tilde{\bm{u}}\in\tilde{\mathcal{U}}$ if and only if $\tilde{\bm{u}}$ satisfies
@@
for all topics $k=1,\ldots,K$ and all subsets $S\subseteq\{1,\ldots,K\}$. We also note that $\overline{\tilde{\mathcal{U}}}$ is bounded because the points it comprises satisfy $\|\bm{u}\|_{d,k}<c$ (see Footnote 8). Therefore, $\tilde{\mathcal{U}}$ is the interior of a bounded convex polytope, and its vertices satisfy a linear system comprised of equations of the form
@@
(64) $\bm{1}_{S}]^{k}=\begin{cases}-1&k\in S\,,\\ 1&\textrm{otherwise}\,.\end{cases}$
@@
Suppose $\bm{u}=\bm{1}_{S}\odot\bm{v}$, where $S\subseteq\{1,\ldots,K\}$ and $\bm{v}$ is a vertex of $\mathcal{U}$. By Lemma 3.9, we must have $\bm{v}\in\mathcal{V}_{s}$ for $s\in\{0,1,\ldots,K\}$. We claim that $\bm{u}$ is a vertex if and only if $s>0$. Suppose $s>0$. We note that $\bm{v}$ must be the solution of a linear system with $s$ equations of the form (52) and KsK-s equations of the form (53). Without loss of generality, we can assume that $\bm{v}$ is a solution of the linear system consisting of (52) for $k=1,\ldots,s$ and (53) for $k=s+1,\ldots,K$. This implies that $\bm{u}$ is a solution to the system consisting of equations (63) for $k=1,\ldots,s$ and (53) for $k=s+1,\ldots,K$. However, it is readily checked that this system is equivalent to the system consisting of equations (63) for $k=1,\ldots,s$ and equations
@@
In Fig. 2, we show the region of receptiveness in 2D for a fixed value of $c$ and various values of $\omega$. When $\omega=1$ (i.e., as in the baseline BCMs), the region of receptiveness is a square. As we decrease $\omega$, the area of the region of receptiveness increases; a node has a larger region of opinion space that a neighbor can be in for them to be mutually receptive on all topics.
@@
For the case $K=3$, the region of receptiveness for the opinion vector $\bm{x}$ consists of a polygon with 26 vertices of the form $\bm{x}+\bm{u}$. The $s=3$ case yields 8 vertices, where each entry of $\bm{u}$ is $c$ or c-c. The $s=2$ case yields 12 vertices, where one entry of $\bm{u}$ is 0 and the remaining two entries are ${3c}/({\omega+2})$ or ${3c}/({\omega+2})$. For the case $s=1$, it yields 6 vertices, where two entries of $\bm{u}$ are 0 and the remaining entry is ${3c}/({2\omega+1})$ or ${3c}/({2\omega+1})$.
@@ 4 Details of our numerical simulations
In this section, we discuss the setup for our numerical simulations of our topic-weighted BCMs. For our simulations, we consider opinions with $K=2$ topics that lie in the region $0,1]\times[0,1]$. For this bounded opinion space, we correspondingly examine confidence bounds999 When $c=0$, adjacent nodes $i$ and $j$ are never receptive to each other on any topic $k$. When $c=1$, adjacent nodes $i$ and $j$ are always receptive to each other on any topic $k$. We do not examine these values of $c$. $c\in(0,1)$.
@@ 4.2 Characterizing opinion fragmentation in our simulations
Researchers have introduced various notions of opinion polarization and fragmentation [11] and have developed multiple methods to quantify them [11, 17, 1]. Intuitively, one simple way to quantify opinion fragmentation is to look at the number of opinion clusters, with a larger number of opinion clusters indicating more opinion fragmentation. However, this does not take into account the cluster sizes (i.e., the number of nodes in a cluster). Suppose that there are two opinion clusters. If the two opinion clusters are the same size, then one can view the opinions in the system as being more fragmented than if one opinion cluster has a large majority of the nodes and the other opinion cluster has a small minority. In our numerical simulations, there can be considerable variation in the sizes of the opinion clusters. Therefore, it is important to consider both the numbers of opinion clusters and their sizes to quantify opinion fragmentation.
To quantify opinion fragmentation in our numerical simulations, we use a modification of the order parameter $Q$ defined by Wang et al. [55]. We calculate [^11] Wang et al. [55] used the inequality $|x_{i}-x_{j}|\leq c$ in their calculation of $Q$ for the standard HK model. However, we use a strict inequality to be consistent with the strict inequality in the update rules (7) and (8). the order parameter $Q$ as
@@
where $\bm{1}_{d_{\mathrm{max}}(\bm{x}_{i}(t),\bm{x}_{j}(t))<c}$ is an indicator function that equals 1 when $d_{\mathrm{max}}(\bm{x}_{i}(t),\bm{x}_{j}(t))<c$ and equals 0 otherwise. The order parameter $Q$ indicates the fraction of pairs of nodes that are receptive to each other on all topics. When $Q=1$, all nodes are receptive to each other on all topics, and we say that nodes reach a consensus state in the infinite-time limit, which is an opinion state in which all opinion vectors are equal. When $Q<1$, we say that the nodes are in a fragmented state. We say that smaller values of $Q$ indicate more opinion fragmentation. For a complete graph, the order parameter gives the fraction of edges $(i,j)\in E$ that are in the effective graph. In a graph that is not complete, some nodes are not adjacent to each other, so the order parameter is more difficult to relate to the effective graph.
@@
For each of our simulations, we characterize opinions by calculating the final order parameter $Q(T_{f})$ at the time at which our simulation reaches the stopping criterion. Let $Q^{*}$ denote the limit order parameter, which is the order parameter (see (70)) calculated with the limit opinion vectors $\bm{x}_{i}^{*}$ for each node $i$. If $Q(T_{f})=1$, then there is a single final opinion cluster and all pairs of adjacent nodes are receptive to each other on all topics. In this scenario, nodes will compromise their opinions over time and eventually reach a consensus state, yielding a limit order parameter $Q^{*}=1$. If $Q(T_{f})<1$, then there are at least two final opinion clusters. In this scenario, the final opinion clusters need not be the same as the opinion clusters of $G_{\mathrm{eff}}^{*}$, and it is possible that $Q(T_{f})\neq Q^{*}$. However, for a small tolerance value, $Q(T_{f})$ is a reasonable approximation of $Q^{*}$.
@@ 4.3 Initial opinion distributions
if $x^{1},x^{2}\in[0,1]$ and 0 otherwise, where $\delta$ is the Dirac delta function. We show this wedge distribution in Fig. 3. The marginal opinion distributions for $x^{1}$ and $x^{2}$ (the opinions on topics 1 and 2, respectively) are uniform on $[0,1]$, which is identical to the independent case. However, $x^{1}$ and $x^{2}$ are highly correlated. To sample from this wedge distribution $f(x^{1},x^{2})$, we draw $x^{1}$ first uniformly at random from $[0,1]$ and then compute $x^{2}$ using $x^{2}=1-|2x^{1}-1|$.
@@
if $x^{1},x^{2}\in[0,1]$ and 0 otherwise. Here, $x^{1}$ and $x^{2}$ are the opinions on topics 1 and 2, respectively, $\sigma$ is the standard deviation of the opinions on topics 1 and 2, and $\rho$ is the Pearson correlation between the opinions on topics 1 and 2. In particular, we consider two truncated Gaussian distributions in our simulations: (1) $G(0.22,0)$, where $x^{1}$ and $x^{2}$ have the same variance and are independent; and (2) $G(0.22,0.8)$, where $x^{1}$ and $x^{2}$ have the same variance but are correlated. To draw opinions randomly from a truncated Gaussian distribution, we first draw randomly from a non-truncated Gaussian distribution. If the sample does not lie in $0,1]\times[0,1]$, we reject it and resample using the same procedure. For our choice of $\sigma=0.22$, we expect that 95% of the random samples lie within the opinion space $0,1]\times[0,1]$, which makes this simple rejection sampling method effective.